OpenStudy (anonymous):

The leg of a right triangle is 4 units and the hypotenuse is 9 units. What is the length, in units, of the other leg of the triangle?

OpenStudy (campbell_st):

using pythagoras theorem a and b are shorter legs and c is the hypotenuse the theorem says \[c^2 = a^2 + b^2\] you can rewrite this to find a shorter leg \[a^2 = c^2 - b^2\] you need to let c = 9 and b = 4 then solve for a

OpenStudy (anonymous):

okay, what about this one? The legs of a right triangle are 6 units and 7 units. What is the length of the hypotenuse?

OpenStudy (campbell_st):

so use the 1st version of the formula \[c^2 = a^2 + b^2\] a and b are shorter legs, c is the hypotenuse

OpenStudy (anonymous):

so its the same thing basically?